3+5+6=151872,5+5+6=253094,5+6+7=303585,5+5+3=251573 than What is 9+4+7=

Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(53756)

.
Considered as is, without explanations (which MUST follow) it is TRASH.


Answer by Edwin McCravy(20077)


The first part of the pattern is easy, if you separate
the digits on the right: 

3+5+6=15 18 72
5+5+6=25 30 94
5+6+7=30 35 85
5+5+3=25 15 73

The number formed by the 1st and 2nd digits on
the right is the product of the first two numbers 
on the left.  
The number formed by the 3rd and 4th digits on
the right is the product of the first and third 
numbers on the left.

So we know eveything but the last two digits:

9+4+7=3663??

We only have to figure out a pattern to go
from the 4 numbers 

3,5,6,15,18 to 72,
5,5,6,25,30 to 94,
5,6,7,30,35 to 85,
5,5,3,25,15 to 73.

The sum of the last two numbers are

15+18 = 33, not close to 72, but noticeably close to its reverse 27
25+30 = 55, not close to 94, but noticeably close to its reverse 49
30+35 = 65, not close to 85, but noticeably close to its reverse 58
25+15 = 40, not close to 73, but noticeably close to its reverse 37

Let's see how close they are to those reverses?

33 is 6 more than 27
55 is 6 more than 49
65 is 7 more than 58
40 is 3 more than 37

Now we notice that 6,6,7,3 are the third numbers.

So the pattern is

1st 4 digits on the right = the reverse of: 
the sum of (the product of the 1st and 2nd on left) 
and (the product of the 3rd and 4th on left),
decreased by the 3rd digit.

So for 9+4+7=3663??

We add 36+63, get 99, then we subtract the 3rd digit
on the left, 7, and get 92.  Then we turn it backward
as 27.
 
Edwin